OpenStudy (anonymous):

cos theta= - square root 2 divided by 2

OpenStudy (mertsj):

\[\cos \theta= -\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

yeah I need to solve the equation (0<theta<2pi)

OpenStudy (mertsj):

We see that the cosine is negative. In what quadrants is the cosine negative?

OpenStudy (anonymous):

2 and 3

OpenStudy (mertsj):

And on your unit circle, what angle has sqrt2/2 for its cosine?

OpenStudy (anonymous):

-45

OpenStudy (mertsj):

45 does have that cosine ut we said we would be in quadrants II and III where the cosine is negative. Also you said your answer had to be between 0 and 2 pi so I think you should give the answer in radians.

OpenStudy (mertsj):

How many radians is 45 degrees?

OpenStudy (anonymous):

45 degrees is pi/4

OpenStudy (mertsj):

So...in quadrants II and III:

OpenStudy (mertsj):